Question text


An aeroplane flies at a ground velocity (i.e. velocity relative to



the ground) of 300 km/h N 30o W, in a wind blowing at a velocity



of 50 km/h N 20o E. What is the velocity (speed and direction)



of the plane relative to the ground? (Use a calculator and round the speed



to the nearest km/h, and the corresponding angle to the nearest degree.)


1
Expert's answer
2022-06-09T14:05:38-0400

a.


"\vec{v}=300\cos120\degree \vec{i}+300\sin120\degree\vec{j}"

"\vec{u}=50\cos70\degree \vec{i}+50\sin70\degree\vec{j}"


"\vec{v}_{res}=\vec{v}+\vec{u}"


"=(300\cos120\degree+50\cos70\degree )\vec{i}"



"+(300\sin120\degree+50\sin70\degree)\vec{j}"

"\approx-132.9\vec{i}+306.8\vec{j}"


"|\vec{v}_{res}|\approx\sqrt{(-132.9)^2+(306.8)^2}\approx334.35(km/h)"



b.


"\tan \theta=\dfrac{306.8}{-132.9}=-2.3085"




"\theta=\tan^{-1}(-2.3085)+180\degree\approx113.42\degree"



An aeroplane travels in the direction N "23.42\degree"W at "334.35" km/h.